From GlueXWiki
Jump to: navigation, search

December 18, 2014 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. CDC update
    amp vs n
    time vs n
    • Early straws are from HVB C8: roc 28, slot 15, channels 48-71; ring 8 straws 53-60, ring 9 straws 62-69 and ring straws 61-68
  2. FDC update (Lubomir, Vlad)
  3. Other: electronics, engineering


Participants: Naomi (CMU), Beni (Switzerland), Mike, Dave, Curtis, Chris, Simon, Vlad, and Lubomir (JLab).

CDC update

- Mike: trying to explain the factor of 4 higher gains (both with cosmics and with beam) now as compared to the tests at CMU. Mike found that the CO2 gas factor is now 0.77 instead of the recommended 0.74 but this could not explain the above factor of 4. HV change (Naomi) is also very unlikely since you need at least 150V for that. Most likely the factor of 4 is coming from the electronics/analysis and Naomi is now setting up the DAQ with the prototype straws to figure this out. Beni: when the CDC was in the EEL126 it was tested only with a scope by Anatoly, but not with the DAQ. Chris will try to figure out which pre-amps have been used at CMU.

- Naomi discussed some plots (linked above) showing that the signals from 24 channels (one connector) are coming earlier than the rest by about 30-40ns, as seen from the fADC125 data. NO idea what's the reason for that.

- According to Beni, for the Holiday break we can lower the CDC flow to about 0.8l/min (from 1.3l/min now). For the FDC we can go down to 80ccpm per package (from 100ccpm).

FDC update

- We had 6 hours of very productive running (runs 2241-2244 total of ~2.5M events or 1.5M good tracks) with no field and FCAL&TOF trigger. We used 2x10^-5 radiator with 100nA, above this the FDC would trip from time to time at 5uA threshold. The tracks projected on the TOF are shown above: few more active TOF paddles on the right are visible, also the inner +/-12cm FCAL blocks were excluded from the trigger.

- Second plot above shows the drift time from the first and fourth packages. One can see a time difference of ~8.5ns, while one would expect ~5.3ns (156cm between packages) for relativistic particles. If true we are dealing with low energy stuff, i.e. 100-200MeV pions??? Another issue in the plot: 5ns structure in the timing??

- Vlad showed plots from the latest runs indicating that we have much more hits per event with the new FCAL&TOF trigger compared with the previous runs with FCAL or BCAL trigger (first plot). Also he showed that for the gain calibration most likely one needs at least 60k hits per plane, the gain coefficient distributions are wider with lower statistics.

High luminosity studies

- The contribution of the active target to the current in the CDC is more than a factor of five as compared at 1x10^-4 radiator and 100nA, 1200A. The pedestal sagging was correspondingly lower. At the maximum luminosity at 3x10^-4 and 150nA the sagging was an most 30 fADC units. At this luminosity, in the FDC currents didn't exceed 5uA per HV channel and we observed some minor pile-up of a few fADC units on the most inner long strips. All this means that when having vacuum in front of the magnet there will be no problems to operate the drift chambers at the GLueX nominal "low luminosity".


- The FDC chiller will be shipped for repair first week of January. We expect it back at the end of February. Curtis: according to Eugene the spring run for Hall D will start at the beginning of April.

- Naomi asked about the plans for reducing the noise from the CAEN HV modules: Chris and Nick are planing to work on re-routing the HV cables in the CDC, Fernando discussed with Eugene the option for installation of filters on the back of the modules.